Grade 11Physical Chemistry

A 150ml of solution of i2 is divided into two unequal parts. I part reacts with hypo solution in acidic medium. 15ml of 0.4M hypo was consumed. II part was added with 100ml of 0.3M NaoH solution residual base required 10ml of 0.3M H2SO4 solution for complete neutralization. What was initial concentration of I2

Aditya kulkarni
9 Years agoGrade 11

1 Answer

Piyush Kumar Behera
9 Years ago
Na2S2O3+I2=Na2S4O6+NaI
Milli moles of Na2S2Oconsumed=6
millimoles of I2 reacted with naoh=12
total millimoles of Iconsumed in the reaction=15
initial concentration=15/150=0.1M